OpenStudy (anonymous):

Find an equation of the tangent line to the given curve at the specified point. y = ex x , (1, e)

OpenStudy (aum):

Is y = \(\large e^xx\) ?

OpenStudy (anonymous):

is that at the point (1,e)??

OpenStudy (aum):

It is not cleat what y is and so I am asking if the function is e raised to x times x as shown in my first reply.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

my bad

OpenStudy (aum):

To find the equation of the tangent we need the slope of the tangent. Find the derivative dy/dx first. What do you get?

OpenStudy (anonymous):

e^x is the slope

OpenStudy (aum):

Could you post a screenshot of the question?

OpenStudy (anonymous):

and this question too please

OpenStudy (zzr0ck3r):

what is the derivative of \(\frac{e^x}{x}\) ?

OpenStudy (anonymous):

its (1/x - 1/x^2)e^x

OpenStudy (aum):

The derivative can be simplified to: \(\large y' = \frac{e^x(x-1)}{x^2} \) Evaluate the derivative at x = 1 y' = 0. Therefore, the slope of the tangent to the curve at (1,e) is zero. That means the tangent is a horizontal line parallel to the x-axis since the slope is zero. Horizontal lines have the equation y = constant. And since this tangent passes through the point (1,e) where the y-value is 'e', the equation of the tangent is y = e.

OpenStudy (anonymous):

that worked, thanks

OpenStudy (aum):

For the second question, I will do one and you can try the rest. a) h(x) = 5f(x) - 4g(x) h'(2) = ? h(x) = 5f(x) - 4g(x) h'(x) = 5f'(x) - 4g'(x) h'(2) = 5f'(2) - 4g'(2) = 5*(-3) - 4*1 = -15 - 4 = -19